Using Appendix 2 in your textbook, what is the value for the standard change in Gibbs free energy (in kJ) at 25oC of the reaction: C2H4(g) + H2(g) --> C2H6(g)?
Answer #1

Standard Gibbs free energy of ethane,

\Delta G_{f}^{o}(C_{2}H_{6}) = - 32.9 kJ/mol

Standard Gibbs free energy of ethylene,

\Delta G_{f}^{o}(C_{2}H_{4}) = 68.2 kJ/mol

Standard Gibbs free energy of hydrogen,

\Delta G_{f}^{o}(H_{2}) = 0 kJ/mol

\Delta G_{rxn}^{o} = \sum \Delta G_{f}^{o}(products) - \sum \Delta G_{f}^{o}(reactants)

                 = - 32.9 - (68.2 + 0)

                 = - 101.1 kJ/mol

Therefore, the standard change in Gibbs free energy for the reaction = - 101.1 kJ/mol